(11%) Problem 9: A heart defibrillator being used on a patient has an RC time constant of 10.5 ms due to the resistance of thall please
Answer #1

Time constant =Rc 10.5x10 3 = RX7.5x10-6 R = 1.4 kn v=Voethe 0.60=146-4/10-5 It=33.07 ms​​​​​​